kristen00001:

Find the slope of the line containing the points(-3,1) and (3,4).

CripQUEZZ:

ok do u know the formula?

kristen00001:

nope

CripQUEZZ:

on the formula is y2 - y1/ x2 - x1

CripQUEZZ:

ok*

kristen00001:

ok ty

kristen00001:

but i still dont know the answer

RodWav3:

i need the formula

CripQUEZZ:

yeah so 4 - 1/ 3 - (-3)

RodWav3:

and try step by step

kristen00001:

they game a multiple choice

kristen00001:

gave*

RodWav3:

oh

CripQUEZZ:

i but the -3 in parentheses bc its a negative

RodWav3:

what is the choices

CripQUEZZ:

so solve this

@cripquezz wrote:
yeah so 4 - 1/ 3 - (-3)
solve this

kristen00001:

m= -1/2 m=2 m=-2 m=1/2

kristen00001:

okk

CripQUEZZ:

can u show a ss of the answers?

kristen00001:

i got 6 2/3

CripQUEZZ:

the answer would be 3/6 and simplified it would be 1/2

kristen00001:

ok tysmmm

CripQUEZZ:

ywww